Show that Re (iz) = −Im (z) and Im (iz) = Re (z). Evaluate Re (1/z) and Im (1/z) if



z = x + iy and z 6= 0.


1
Expert's answer
2021-12-20T16:48:26-0500

"Re(z)=x,Im(z)=y"

"iz=-y+ix"

"Re(iz)=-y=Im(z)"

"Im(iz)=x=Re(z)"


"1/z=\frac{x-iy}{(x+iy)(x-iy)}=\frac{x-iy}{x^2+y^2}"


"Re(1/z)=\frac{x}{x^2+y^2}"


"Im(1/z)=-\frac{y}{x^2+y^2}"